Question 1


Stonewall receives ¢250 per year in simple interest from an amount of money he invested in ADB, Barclays and GCB. Suppose ADB pays an interest of 2%, Barclays pays an interest of 4% and GCB pays an interest of 5% per annum and an amount of ¢350 more was invested in Barclays than the amount invested in ADB and GCB combined. Also, the amount invested in Barclays is 2 times the amount invested in GCB.


a) Write down the three linear equations and represent them in the matrix form ?? = ?.


b) Find the amount of money Stonewall invested in ADB, Barclays and GCB using Matrix inversion.






1
Expert's answer
2021-07-06T03:24:12-0400

Lets x - invested in ADB,

y - invested inBarclays ,

z - invested in GCB.

"0.02x+0.04y+0.05z=250"

"y=350+(x+z)"

"y=2z"

We have a system

"0.02x+0.04y+0.05z=250\\
-x+y-z=350\\
y-2z=0"

"A=\begin{pmatrix}
 0.02 & 0.04&0.05 \\
 -1& 1&-1\\
0&1&-2
\end{pmatrix}, X=\begin{pmatrix}
 x \\
 y\\
z
\end{pmatrix}, \\B=\begin{pmatrix}
 250\\
 350\\
0
\end{pmatrix}"

"AX=B\\
X=A^{-1}\cdot B\\
A^{-1}:"


"\begin{pmatrix}
 0.02 & 0.04&0.05&|1&0&0 \\
 -1& 1&-1&|0&1&0\\
0&1&-2&|0&0&1
\end{pmatrix}\leftrightarrow"

Ir"\leftrightarrow"IIr

"\begin{pmatrix}
 -1&1&-1&|0&1&0 \\
 0.02&0.04&0.05&|1&0&0\\
0&1&-2&|0&0&1
\end{pmatrix}\leftrightarrow"

IIr+Ir(0.02)

Ir(-1)

"\begin{pmatrix}
 1&-1&1&|0&-1&0 \\
 0&0.06&0.03&|1&0.02&0\\
0&1&-2&|0&0&1
\end{pmatrix}\leftrightarrow"

IIr"\leftrightarrow" IIIr

"\begin{pmatrix}
 1&-1&1&|0&-1&0 \\
 0&1&-2&|0&0&1\\
0&0.06&0.03&|1&0.02&0\\
\end{pmatrix}\leftrightarrow"

IIIr+IIr(-0.06)

"\begin{pmatrix}
 1&-1&1&|0&-1&0 \\
 0&1&-2&|0&0&1\\
0&0&0.15&|1&0.02&-0.06\\
\end{pmatrix}\leftrightarrow"

IIIr("\frac{20}{3}")

"\begin{pmatrix}
 1&-1&1&|0&-1&0 \\
 0&1&-2&|0&0&1\\
0&0&1&|\frac{20}{3}&\frac{2}{15}&-\frac{2}{5}\\
\end{pmatrix}\leftrightarrow"

IIr+IIIr(2)

Ir+IIIr(-1)

"\begin{pmatrix}
 1&-1&0&|-\frac{20}{3}&-\frac{17}{15}&\frac{2}{5} \\
 0&1&0&|\frac{40}{3}&\frac{4}{15}&\frac{1}{5}\\
0&0&1&|\frac{20}{3}&\frac{2}{15}&-\frac{2}{5}\\
\end{pmatrix}\leftrightarrow"

Ir+IIr

"\begin{pmatrix}
 1&0&0&|\frac{20}{3}&-\frac{13}{15}&\frac{3}{5} \\
 0&1&0&|\frac{40}{3}&\frac{4}{15}&\frac{1}{5}\\
0&0&1&|\frac{20}{3}&\frac{2}{15}&-\frac{2}{5}\\
\end{pmatrix}"


"A^{-1}=\begin{pmatrix}
 \frac{20}{3}&-\frac{13}{15}&\frac{3}{5} \\
 \frac{40}{3}&\frac{4}{15}&\frac{1}{5}\\
\frac{20}{3}&\frac{2}{15}&-\frac{2}{5}\\
\end{pmatrix}"

"X=\begin{pmatrix}
 \frac{20}{3}&-\frac{13}{15}&\frac{3}{5} \\
 \frac{40}{3}&\frac{4}{15}&\frac{1}{5}\\
\frac{20}{3}&\frac{2}{15}&-\frac{2}{5}\\
\end{pmatrix}\cdot\begin{pmatrix}
 250\\
 350\\
0
\end{pmatrix}=\\
=\begin{pmatrix}
 \frac{4090}{3}\\
 \frac{10280}{3}\\
\frac{5140}{3}
\end{pmatrix}"

x=1363"\frac{1}{3}"

y=3426"\frac{2}{3}"

z=1713"\frac{1}{3}"